OpenStudy (anonymous):

At a fruit stand,apples can be bought for $0.15 each and pears for $0.20 each.At these rates, a bag of apples and pears was bought for $3.80.If the bag contained exactly 21 pieces of fruit,how many were pears ? (note :How do you solve this problem ?The answer is supposed to be 13)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

don't you hate word problems?

OpenStudy (anonymous):

you can write \[x+y=21\] and \[.15x+.20y=3.80\] if you like

OpenStudy (anonymous):

second one would be easier if you write it using pennies instead of dollars \[x+y=21\\ 15x+20y=380\] then solve that one

OpenStudy (anonymous):

the first is the same as \(y=21-x\) so you can even solve \[15x+20(21-x)=380\] for \(x\)

OpenStudy (anonymous):

let me know what you get, or if you are sill confused i wrote the equations without much explanation

OpenStudy (anonymous):

@AcademicGurusInc are you writing a reply or posting a link to a commercial site? the first is helpful, please chime in the second is not

OpenStudy (anonymous):

Thank you so much AcademicGurusInc and satellite 73 for your help ! I understand know how to do it .

OpenStudy (anonymous):

Hi misty1212,yes word problems are awful

OpenStudy (misty1212):

lol i hate them

OpenStudy (academicgurusinc):

You're welcome :)